Fifty matched pairs of​ magnitude/depth measurements randomly selected from​ 10,594 earthquakes recorded in one year from a location in southern California. Find the best predicted depth of an earthquake with a magnitude of 1.8. Use a significance level of 0.05 from table = .279.

Mag   Depth
2.61   13.5
0.78   11.9
2.63   19.4
1.51   13.2
1.73   9.2
0.22   11.7
2.22   16.9
0.79   3.9
0.02   8.7
2.66   6.8
1.25   18.8
0.98   7.7
1.95   19.4
2.63   2.1
2.41   11.5
1.77   3.8
1.82   14.7
0.17   19.8
0.36   3.5
0.94   19.5
1.87   10.7
0.99   10.7
1.62   6.2
2.44   4.9
0.52   11.7
2.74   12.1
0.79   14.5
2.68   17.4
2.21   5.7
2.86   15.7
0.29   3.8
0.48   9.3
2.04   4.4
2.62   11.7
0.83   10.1
1.47   8.9
2.69   11.5
0.21   16.4
1.15   10.1
0.43   16.8
2.46   13.8
2.49   17.5
1.17   13.5
2.99   18.1
0.75   11.7
1.45   10.9
2.04   12.8
2.12   15.2
1.51   9.3
1.67   11.5

The regression equation is y (with caret) ___ + ___x. ​(Round to one decimal place as​ needed.)

Given that the earthquake has a magnitude of 1.8​, the best predicted depth is ___km. ​(Round to one decimal place as​ needed.)
Answer #1

applying regression on above data:

71 1 3 0. 59 2 6 0 27, 60 8 23. 51 15 59 91 66 61 0 29 45 0 8 32 35 16 6 5 52 2 7 0 2 2 26 4 34 3 41 0 1150 79978 9715477 471 2. 0 2 1 1 0 2 0 0 2 1 0 1 2 2 1 1 0 0 0 1 0 1 2 0 2 0 2 2 2 0 0 2 2 0 1 2 0 1 0 2 2 1 2 0 1 2 2 1

regression equation is y (with caret) =10.8+0.6*x

best predicted depth is =10.8+0.6*1.8 =11.9 km